# The torus plateau for the high-dimensional Ising model

### v1

@@ 3. $(iii)$
By definition, $\overline{\Gamma}_{\mathbf{m}}$ is determined by $\Gamma_{\mathbf{m}}$ and the ordering $\prec$, and similarly $\Gamma_{\mathbf{m}}$ can be determined from $\overline{\Gamma}_{\mathbf{m}}$ by selecting odd edges in $\overline{\Gamma}_{\mathbf{m}}$ according to $\prec$. Also, the current $\mathbf{m}\setminus\overline{\Gamma}_{\mathbf{m}}$, defined to be the restriction of $\mathbf{m}$ to the complement $\overline{\Gamma}_{\mathbf{m}}^{c}$, is sourceless.
@@
Since $\partial\mathbf{m}=\{y,y^{\prime}\}$ now, we can form the backbone $\Gamma_{\mathbf{m}}$ of $\mathbf{m}$ from $y^{\prime}$ to $y$. Let $\overline{\Gamma}_{\mathbf{m}}$ be the set of edges explored during the formation of $\Gamma_{\mathbf{m}}$. Let $\mathcal{C}_{\mathbf{m}\setminus\overline{\Gamma}_{\mathbf{m}}+\hat{\mathbf{n}}}(0)$ denote the result of the exploration process of Section 3.2 performed on the current $\mathbf{m}\setminus\overline{\Gamma}_{\mathbf{m}}+\hat{\mathbf{n}}$ (on the graph ${\Lambda_{R}}\setminus\overline{\Gamma}_{\mathbf{m}}$). We claim that
@@
Indeed, if $\mathcal{C}_{\mathbf{m}\setminus\overline{\Gamma}_{\mathbf{m}}+\hat{\mathbf{n}}}(0)=\mathcal{C}(0)$, then (3.33) holds with $z=y^{\prime}$. If $\mathcal{C}_{\mathbf{m}\setminus\overline{\Gamma}_{\mathbf{m}}+\hat{\mathbf{n}}}(0)\neq\mathcal{C}(0)$, then there is a first edge $zv\in\overline{\Gamma}_{\mathbf{m}}$ used in the construction of $\mathcal{C}(0)$ that makes the two exploration processes different. For this edge, we must have $z\in\mathcal{C}_{\mathbf{m}\setminus\overline{\Gamma}_{\mathbf{m}}+\hat{\mathbf{n}}}(0)$ and $\mathbf{m}_{zv}>0$. Since one of $z,v$ must be in $\Gamma_{\mathbf{m}}$, we see that $z\xleftrightarrow{\overline{\Gamma}_{\mathbf{m}}}y^{\prime}$ also. This proves (3.33).
@@
Inserting this yields
@@ Appendix A Torus convolution estimates
We present two elementary convolution lemmas. The first is a special case of [34, Lemma 3.4]. We include the simple proof for completeness, and as an opportunity to correct an omission in the proof of [34, Lemma 3.4] (repeated in [33, Lemma 4.2]) which however does not change the conclusion of those lemmas.
@@
The next lemma is a minor extension of [34, Lemma 3.6]. The definition of $T_{\beta}^{(r)}(x)$ is in (1.24), $\star$ denotes convolution on $\mathbb{T}_{r}$ and * denotes convolution on $\mathbb{Z}^{d}$.
@@
where $Z^{A}=\sum_{\partial\mathbf{n}=A}w(\mathbf{n})$ is the normalisation constant. Given two vertex sets $A_{1},A_{2}$, we define
@@
We write $\mathcal{P}_{2}=\mathcal{P}_{2}({\mathbb{T}_{r}})$ for the set of all unordered pairs of vertices in $\mathbb{T}_{r}$. Let $\mathbf{C}(u)$ denote the vertex cluster of $u\in\mathbb{T}_{r}$ in the percolation configuration constructed from a double current $\mathbf{n}_{1}+\mathbf{n}_{2}$, and let $\mathbf{C}^{b}(u)$ denote the set of pairs $\{s,t\}\in\mathcal{P}_{2}$ such that either $s$ or $t$ belongs to $\mathbf{C}(0)$. Given $\theta\in[0,1]$ and $E\subset\mathcal{P}_{2}$, we introduce the (random) interaction $(J^{E}_{u,v}(\theta))_{\{u,v\}\in\mathcal{P}_{2}}$ defined by
$J^{E}_{u,v}(\theta)=\begin{cases}\theta J_{u,v}&(\{u,v\}\in\mathbf{C}^{b}(0)\cap E)\\ J_{u,v}&(\{u,v\}\not\in\mathbf{C}^{b}(0)\cap E),\end{cases}$ (B.9)
and we write $\langle\cdot\rangle_{\theta}$ for the Ising expectation with the random interaction $J^{E}(\theta)$. Given a subset $G\subset\mathcal{P}_{2}$, we regard $G$ as a graph whose edges consist of the elements of $G$ and whose vertices are the vertices contained in the elements of $G$, and we write $\langle\cdot\rangle^{G}$ for the Ising expectation on the graph $G$ with interaction $J$. By definition, $\langle\cdot\rangle_{1}=\langle\cdot\rangle$ and $\langle\cdot\rangle_{0}=\langle\cdot\rangle^{(\mathbf{C}^{b}(0)\cap E)^{c}}$. According to our convention, the graph $(\mathbf{C}^{b}(0)\cap E)^{c}$ consists of the edges not in $\mathbf{C}^{b}(0)\cap E$, together with the vertices comprising those edges.
@@
$T_{2}(x,y,z,E)$ $=\beta\sum_{\{u,v\}\in E}J_{u,v}\langle\sigma_{0}\sigma_{x}\rangle\mathbf{E}^{0x,\varnothing}\bigg{[}\mathds{1}[u\textup{ or }v\in\mathbf{C}(0)]\int_{0}^{1}\theta\;\partial_{\theta}\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle_{\theta}\mathrm{d}\theta\bigg{]}.$ (B.11)
@@
This can be achieved by the classical manoeuvre of peeling off the cluster of 00, by conditioning on $\mathbf{C}^{b}(0)$:
$Z^{0x,yz}[\mathbf{C}(0)\cap\mathbf{C}(y)=\varnothing]$ $=\sum_{H\subset\mathcal{P}_{2}}Z^{0x,\varnothing}_{H}[\mathbf{C}^{b}(0)=H]Z^{\varnothing,yz}_{H^{c}}$
$=\sum_{H\subset\mathcal{P}_{2}}Z^{0x,\varnothing}[\mathbf{C}^{b}(0)=H]\langle\sigma_{y}\sigma_{z}\rangle^{H^{c}}=Z^{0x,\varnothing}\mathbf{E}^{0x,\varnothing}[\langle\sigma_{y}\sigma_{z}\rangle^{\mathbf{C}^{b}(0)^{c}}].$ (B.16)
@@
$\langle\sigma_{y}\sigma_{z}\rangle^{\mathbf{C}^{b}(0)^{c}}\leq\langle\sigma_{y}\sigma_{z}\rangle^{{(\mathbf{C}^{b}(0)\cap E)^{c}}}=\langle\sigma_{y}\sigma_{z}\rangle_{0}.$ (B.17)
@@
=$=\sum_{\{u,v\}\in E}\beta J_{u,v}\mathds{1}[u\textup{ or }v\in\mathbf{C}(0)]\int_{0}^{1}\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle_{\theta_{1}}\mathrm{d}\theta_{1}.$ (B.20)
@@
$=\frac{\beta}{2}\sum_{u,v\in{\mathbb{T}_{r}}}\mathds{1}[\{u,v\}\in E]J_{u,v}\langle\sigma_{0}\sigma_{x}\rangle\mathbf{P}^{0x,\varnothing}[u\textup{ or }v\in\mathbf{C}(0)]\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle.$ (B.24)
@@
$T_{2}(x,y,z,E)=\beta\sum_{\{u,v\}\in E}J_{u,v}\langle\sigma_{0}\sigma_{x}\rangle\mathbf{E}^{0x,\varnothing}\bigg{[}\mathds{1}[u\textup{ or }v\in\mathbf{C}(0)]\int_{0}^{1}\theta\;\partial_{\theta}\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle_{\theta}\mathrm{d}\theta\bigg{]}.$ (B.30)
@@
$\partial_{\theta}\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle_{\theta}=\beta\sum_{\{s,t\}\in E}J_{s,t}\mathds{1}[s\textup{ or }t\in\mathbf{C}(0)]\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v};\sigma_{s}\sigma_{t}\rangle_{\theta},$ (B.31)
@@
$T_{2}(x,y,z,E)$ $\leq\frac{\beta^{2}}{4}\sum_{\begin{subarray}{c}u,v,s,t\in{\mathbb{T}_{r}}\\ \{u,v\}\neq\{s,t\}\end{subarray}}\mathds{1}[\{u,v\},\{s,t\}\in E]J_{u,v}J_{s,t}\langle\sigma_{0}\sigma_{x}\rangle$
×$\qquad\times\mathbf{E}^{0x,\varnothing}\bigg{[}\mathds{1}[u\text{or }v\in\mathbf{C}(0)]\mathds{1}[s\text{ or }t\in\mathbf{C}(0)]\int_{0}^{1}\theta\,\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v};\sigma_{s}\sigma_{t}\rangle_{\theta}\,\mathrm{d}\theta\bigg{]}.$ (B.33)
@@
We insert the above two inequalities into (B), average over $E=\omega$, and then sum over $x,y,z\in{\mathbb{T}_{r}}$. The sums over $x,y,z$ produce $\chi^{3}$. Then, by translation invariance, we can shift the origin to one of $u_{1},u_{2},t_{1},t_{2}$, and then sum over the former origin to get another factor of $\chi$. The remaining sums are of three kinds and can all be written as convolutions. In terms of the torus convolution defined in (2.9), this produces
@@
where the factor 2 is from (B.34) and the first factor 4 is from $(0\Leftrightarrow x)$ and $(y\Leftrightarrow z)$. We then apply Young’s convolution inequality to see that
$\lVert J\star\tau^{2}\star J\rVert_{1}$ $\leq\lVert J\rVert_{1}^{2}\lVert\tau^{2}\rVert_{1},$ (B.37)
@@ Acknowledgements
We thank Hugo Duminil-Copin, Trishen Gunaratnam, and Alexis Prévost for useful discussions. GS thanks the Section de Mathématiques of the University of Geneva for kind hospitality during a visit when this work was initiated. The work of YL and GS was supported in part by NSERC of Canada. The work of RP was supported by the Swiss National Science Foundation, the NCCR SwissMAP, and the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 757296).

### v2

@@ 3. $(iii)$
By definition, $\overline{\Gamma}_{\mathbf{m}}$ is determined by $\Gamma_{\mathbf{m}}$ and the ordering $\prec$, and similarly $\Gamma_{\mathbf{m}}$ can be determined from $\overline{\Gamma}_{\mathbf{m}}$ by selecting the odd edges in $\overline{\Gamma}_{\mathbf{m}}$ according to $\prec$. Also, the current $\mathbf{m}\setminus\overline{\Gamma}_{\mathbf{m}}$, defined to be the restriction of $\mathbf{m}$ to the complement $\overline{\Gamma}_{\mathbf{m}}^{c}$, is sourceless.
@@
Since the current $\mathbf{m}$ on $\Lambda_{R}$ now has sources $\partial\mathbf{m}=\{y,y^{\prime}\}$, we can explore the backbone $\Gamma_{\mathbf{m}}$ of $\mathbf{m}$ from $y^{\prime}$ to $y$. Let $\overline{\Gamma}_{\mathbf{m}}$ be the set of edges explored during the formation of $\Gamma_{\mathbf{m}}$. Let $\mathcal{C}_{\mathbf{m}\setminus\overline{\Gamma}_{\mathbf{m}}+\hat{\mathbf{n}}}(0)$ denote the result of the exploration process of Section 3.2 performed on the current $\mathbf{m}\setminus\overline{\Gamma}_{\mathbf{m}}+\hat{\mathbf{n}}$ (on the graph ${\Lambda_{R}}\setminus\overline{\Gamma}_{\mathbf{m}}$). We claim that
@@
Indeed, if $\mathcal{C}_{\mathbf{m}\setminus\overline{\Gamma}_{\mathbf{m}}+\hat{\mathbf{n}}}(0)=\mathcal{C}(0)$, then (3.34) holds with $z=y^{\prime}$. If $\mathcal{C}_{\mathbf{m}\setminus\overline{\Gamma}_{\mathbf{m}}+\hat{\mathbf{n}}}(0)\neq\mathcal{C}(0)$, then there is a first edge $zv\in\overline{\Gamma}_{\mathbf{m}}$ used in the construction of $\mathcal{C}(0)$ that makes the two exploration processes different. For this edge, we must have $z\in\mathcal{C}_{\mathbf{m}\setminus\overline{\Gamma}_{\mathbf{m}}+\hat{\mathbf{n}}}(0)$ and $\mathbf{m}_{zv}>0$. Since one of $z,v$ must be in $\Gamma_{\mathbf{m}}$, we see that $z\xleftrightarrow{\overline{\Gamma}_{\mathbf{m}}\>}y^{\prime}$ also. This proves (3.34).
@@
Inserting this into (3.38) yields
@@ Appendix A Torus convolution estimates
We present two elementary convolution lemmas. The first is a special case of [36, Lemma 3.4]. We include the simple proof for completeness, and as an opportunity to correct an omission in the proof of [36, Lemma 3.4] (repeated in [35, Lemma 4.2]) which however does not change the conclusion of those lemmas.
@@
The next lemma is a minor extension of [36, Lemma 3.6]. The definition of $T_{\beta}^{(r)}(x)$ is in (1.24), $\star$ denotes convolution on $\mathbb{T}_{r}$ and * denotes convolution on $\mathbb{Z}^{d}$.
@@
where $Z^{A}=\sum_{\partial\mathbf{n}=A}w(\mathbf{n})$ is the normalisation constant. Given two vertex sets $A_{1},A_{2}$, we define the product measure
@@
We write $\mathcal{P}_{2}=\mathcal{P}_{2}({\mathbb{T}_{r}})$ for the set of all unordered pairs of vertices in $\mathbb{T}_{r}$. Let $\mathbf{C}(u)$ denote the vertex cluster of $u\in\mathbb{T}_{r}$ in the percolation configuration constructed from a double current $\mathbf{n}_{1}+\mathbf{n}_{2}$, and let $\overline{\mathbf{C}}(u)$ denote the set of pairs $\{s,t\}\in\mathcal{P}_{2}$ such that either $s$ or $t$ belongs to $\mathbf{C}(u)$. Given $\theta\in[0,1]$ and $E\subset\mathcal{P}_{2}$, we introduce the (random) interaction $(J^{E}_{u,v}(\theta))_{\{u,v\}\in\mathcal{P}_{2}}$ defined by
$J^{E}_{u,v}(\theta)=\begin{cases}\theta J_{u,v}&(\{u,v\}\in\overline{\mathbf{C}}(0)\cap E)\\ J_{u,v}&(\{u,v\}\not\in\overline{\mathbf{C}}(0)\cap E).\end{cases}$ (B.9)
We write $\langle\cdot\rangle_{\theta}$ for the Ising expectation with the random interaction $J^{E}(\theta)$. Given a subset $G\subset\mathcal{P}_{2}$, we regard $G$ as a graph whose edges consist of the elements of $G$ and whose vertices are the vertices contained in the elements of $G$, and we write $\langle\cdot\rangle^{G}$ for the Ising expectation on the graph $G$ with interaction $J$. By definition, $\langle\cdot\rangle_{1}=\langle\cdot\rangle$ and $\langle\cdot\rangle_{0}=\langle\cdot\rangle^{(\overline{\mathbf{C}}(0)\cap E)^{c}}$. According to our convention, the graph $(\overline{\mathbf{C}}(0)\cap E)^{c}$ consists of the edges not in $\overline{\mathbf{C}}(0)\cap E$, together with the vertices comprising those edges.
@@
$T_{2}(x,y,z,E)$ $=\beta\sum_{\{u,v\}\in E}J_{u,v}\langle\sigma_{0}\sigma_{x}\rangle\mathbf{E}^{0x,\varnothing}\bigg{[}\mathds{1}\{u\textup{ or }v\in\mathbf{C}(0)\}\int_{0}^{1}\theta\;\partial_{\theta}\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle_{\theta}\mathrm{d}\theta\bigg{]}.$ (B.11)
@@
This can be achieved by the classical manoeuvre of peeling off the cluster of 00, by conditioning on $\overline{\mathbf{C}}(0)$:
$Z^{0x,yz}[\mathbf{C}(0)\cap\mathbf{C}(y)=\varnothing]$ $=\sum_{H\subset\mathcal{P}_{2}}Z^{0x,\varnothing}_{H}[\overline{\mathbf{C}}(0)=H]Z^{\varnothing,yz}_{H^{c}}$
$=\sum_{H\subset\mathcal{P}_{2}}Z^{0x,\varnothing}[\overline{\mathbf{C}}(0)=H]\langle\sigma_{y}\sigma_{z}\rangle^{H^{c}}=Z^{0x,\varnothing}\mathbf{E}^{0x,\varnothing}[\langle\sigma_{y}\sigma_{z}\rangle^{\overline{\mathbf{C}}(0)^{c}}].$ (B.16)
@@
$\langle\sigma_{y}\sigma_{z}\rangle^{\overline{\mathbf{C}}(0)^{c}}\leq\langle\sigma_{y}\sigma_{z}\rangle^{{(\overline{\mathbf{C}}(0)\cap E)^{c}}}=\langle\sigma_{y}\sigma_{z}\rangle_{0}.$ (B.17)
@@
=$=\sum_{\{u,v\}\in E}\beta J_{u,v}\mathds{1}\{u\textup{ or }v\in\mathbf{C}(0)\}\int_{0}^{1}\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle_{\theta_{1}}\mathrm{d}\theta_{1}.$ (B.20)
@@
$=\frac{\beta}{2}\sum_{u,v\in{\mathbb{T}_{r}}}\mathds{1}\{\{u,v\}\in E\}J_{u,v}\langle\sigma_{0}\sigma_{x}\rangle\mathbf{P}^{0x,\varnothing}[u\textup{ or }v\in\mathbf{C}(0)]\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle.$ (B.24)
@@
$T_{2}(x,y,z,E)=\beta\sum_{\{u,v\}\in E}J_{u,v}\langle\sigma_{0}\sigma_{x}\rangle\mathbf{E}^{0x,\varnothing}\bigg{[}\mathds{1}\{u\textup{ or }v\in\mathbf{C}(0)\}\int_{0}^{1}\theta\;\partial_{\theta}\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle_{\theta}\mathrm{d}\theta\bigg{]}.$ (B.30)
@@
$\partial_{\theta}\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v}\rangle_{\theta}=\beta\sum_{\{s,t\}\in E}J_{s,t}\mathds{1}\{s\textup{ or }t\in\mathbf{C}(0)\}\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v};\sigma_{s}\sigma_{t}\rangle_{\theta},$ (B.31)
@@
$T_{2}(x,y,z,E)$ $\leq\frac{\beta^{2}}{4}\sum_{\begin{subarray}{c}u,v,s,t\in{\mathbb{T}_{r}}\\ \{u,v\}\neq\{s,t\}\end{subarray}}\mathds{1}\{\{u,v\},\{s,t\}\in E\}J_{u,v}J_{s,t}\langle\sigma_{0}\sigma_{x}\rangle$
×$\qquad\times\mathbf{E}^{0x,\varnothing}\bigg{[}\mathds{1}\{u\text{ or }v\in\mathbf{C}(0)\}\mathds{1}\{s\text{ or }t\in\mathbf{C}(0)\}\int_{0}^{1}\theta\,\langle\sigma_{y}\sigma_{z};\sigma_{u}\sigma_{v};\sigma_{s}\sigma_{t}\rangle_{\theta}\,\mathrm{d}\theta\bigg{]}.$ (B.33)
@@
We insert the above two inequalities into (B), average over $E=\omega$, and then sum over $x,y,z\in{\mathbb{T}_{r}}$. The sums over $x,y,z$ produce $\chi^{3}$. Then, by translation invariance, we can shift the origin to one of $u_{1},u_{2},t_{1},t_{2}$, and then sum over the former origin to get another factor of $\chi$. The remaining sums are of three kinds and can all be written as convolutions. In terms of the torus convolution $\star$ defined in (2.9), this produces
@@
where the factor 2 is from (B.34) and the first factor 4 is from $(0\Leftrightarrow x)$ and $(y\Leftrightarrow z)$. We then apply the Cauchy–Schwarz inequality to see that
$\lVert J\star\tau^{2}\star J\rVert_{1}$ $=\lVert J\rVert_{1}^{2}\lVert\tau^{2}\rVert_{1},$ (B.37)
@@ Acknowledgements
We thank Hugo Duminil-Copin, Trishen Gunaratnam, and Alexis Prévost for useful discussions. We also thank two anonymous referees for useful comments. GS thanks the Section de Mathématiques of the University of Geneva for kind hospitality during a visit when this work was initiated. The work of YL and GS was supported in part by NSERC of Canada. The work of RP was supported by the Swiss National Science Foundation, the NCCR SwissMAP, and the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 757296).